# Achieving Unbiased Multi-Instance Learning via Balanced Fine-Grained Positive-Unlabeled Learning

### v2

@@
Abstract.
In real-world applications, it is often challenging to detect anomalous samples when the anomalous information they contain is extremely limited. In such cases, both macro-level and micro-level detection using multi-instance learning (MIL) encounter significant difficulties. The former struggles because normal and anomalous samples are highly similar and hard to distinguish at the macro level, while the latter is limited by the lack of labels at the micro level. In MIL, micro-level labels are inferred from macro-level labels, which can lead to severe bias. Moreover, the more imbalanced the distribution between normal and anomalous samples, the more pronounced these limitations become. In this study, we observe that the MIL problem can be elegantly transformed into a fine-grained Positive-Unlabeled (PU) learning problem. This transformation allows us to address the imbalance issue in an unbiased manner using a micro-level balancing mechanism. To this end, we propose a novel framework—Balanced Fine-Grained Positive-Unlabeled (BFGPU)—based on rigorous theoretical foundations to address the challenges above. Extensive experiments on both public and real-world datasets demonstrate the effectiveness of BFGPU, which outperforms existing methods, even in extreme scenarios where both macro and micro-level distributions are highly imbalanced. The code is open-sourced at https://github.com/BFGPU/BFGPU.
Positive-Unlabeled Learning, Weakly Supervised Learning
1. Introduction
@@
In real-world applications, there is a strong demand for effective detection of anomalous samples, such as in quality inspection, risk control, and fault diagnosis  (Görnitz et al., 2013; Pang et al., 2023; Ye et al., 2023). Ideally, anomalous samples exhibit significant and easily distinguishable differences from normal samples, making them readily separable based on features or representations.
However, in more realistic scenarios, the differences between anomalous and normal samples are often subtle. Anomalies may manifest as sparse local information, while the rest of the information remains indistinguishable from that of normal samples. As a result, anomalous samples are highly similar to normal ones at a macro level, making them difficult to identify. For example, as illustrated in fig. 1, in a real-world customer service quality inspection task, the goal is to detect instances of impoliteness within extended conversations between customer services and customers. Since the customer service is well-trained, mistakes—if any—typically occur in only a subtle utterance.
@@
Figure 1. This example illustrates the difficulty in distinguishing between normal and anomalous macro samples due to the low proportion of anomalous information within anomalous samples.
Given the difficulty of distinguishing samples at the macro level, a promising approach is to seek solutions at the micro level—by decomposing each macro-level sample into multiple finer-grained components and performing discrimination at this finer resolution. However, such approaches face two major challenges. First, there are no precise labels available at the micro level, so learning must rely solely on macro-level supervision. Second, anomalous samples are already scarce at the macro level  (Chandola et al., 2009; Pang et al., 2023; Perini et al., 2023), and within each anomalous sample, truly anomalous information often constitutes only a small fraction. This results in a severe imbalance along two dimensions at both the macro and the micro level, posing significant challenges for effective learning.
@@
Multiple Instance Learning (MIL) is a paradigm of weakly supervised learning designed for scenarios with imprecise labels  (Waqas et al., 2024; Zhou, 2018; Gao et al., 2022). It structures data into two hierarchical levels: the macro level (referred to as bags) and the micro level (referred to as instances), and utilizes only macro-level labels for training. To address the lack of micro-level supervision, existing MIL methods often heuristically assign bag labels to all instances within the bag  (Ilse et al., 2018; Angelidis and Lapata, 2018; Pang et al., 2023). However, this strategy lacks theoretical grounding and can introduce substantial bias. In terms of data imbalance, the micro-level imbalance—known in MIL literature as the low witness rate problem—has been recognized but remains unresolved  (Carbonneau et al., 2018; Zhang et al., 2022). Meanwhile,macro-level imbalance has been largely overlooked and presents an even greater challenge.
@@
In this paper, instead of directly assigning bag labels to instances, we treat all instances from normal bags as positive instances and all instances from anomalous bags as unlabeled instances. This reformulates the MIL problem as an fine-grained Positive-Unlabeled (PU) learning task. PU learning is a learning paradigm that trains models using only positive and unlabeled data  (Bekker and Davis, 2020). Its core assumption is that both labeled and unlabeled positive instances are drawn from the same underlying distribution, which enables estimation of the distribution of negative instances within the unlabeled set and thus facilitates the separation of positive and negative instances—thereby alleviating the label bias challenge in MIL  (Du Plessis et al., 2015; Kiryo et al., 2017). Building upon this reformulation, we further derive a PU learning loss function based on rigorous theoretical analysis, which simultaneously addresses imbalance at both the macro and micro levels—thereby alleviating the dual imbalance challenge in MIL.
@@
Figure 2. This figure illustrates the different solution paradigms for the detection problem in this paper: coarse-grained positive-negative learning, MIL, and fine-grained positive-unlabeled learning.
We conducted extensive experiments to evaluate the performance of BFGPU across various datasets, including 2 datasets with a maximum imbalance ratio of 5 and 2 datasets with a maximum imbalance ratio of 10. Additionally, we tested the algorithm on a real-world customer service quality inspection dataset. These experiments demonstrated the effectiveness of BFGPU in both diverse datasets and practical applications. Furthermore, we explored scenarios with both macro and micro imbalances, and even in extreme cases where anomalous information is vastly outnumbered, our algorithm consistently performed well.
@@
Our Contributions. 1. We propose a novel perspective that reformulates the MIL problem as a balanced PU learning task, which addresses the issues of bias and imbalance in MIL. 2. We incorporate macro information into the learning process, leading to BFGPU, an algorithm that enables balanced and unbiased macro performance by training at the micro level. 3. We provide a theoretical analysis demonstrating the advantages of BFGPU over macro-level learning and conventional MIL approaches. 4. Extensive experiments validate the superior performance of BFGPU.
2. Related Works
2.1. Positive Unlabeled Learning
@@
PU learning can utilize positive data and unlabeled data to train a classifier that distinguishes between positive and negative data  (Bekker and Davis, 2020). It is reasonable to use the loss of unlabeled data, proportionally subtracted from the loss of positive data, to estimate the loss of negative data. Most PU learning algorithms aim for the model to be unbiased in terms of accuracy, and there are many effective algorithms currently available  (Du Plessis et al., 2015; Kiryo et al., 2017; Chen et al., 2020; Kato et al., 2018; Shi et al., 2018; Sansone et al., 2018; Hou et al., 2018; Hsieh et al., 2019, 2015; Wang et al., 2024). Some PU learning algorithms address the issue of imbalance in the positive-negative ratio within unlabeled data and strive for unbiased average accuracy (AvgAcc)  (Su et al., 2021). Additionally, theoretical research on PU learning has demonstrated its superiority over PN learning under certain conditions  (Niu et al., 2016).
2.2. Multi-Instance Learning
@@
Multi-instance learning (MIL) is a form of weakly supervised learning  (Zhou, 2018) characterized by inexact supervision, where labels are not precise  (Carbonneau et al., 2018). In MIL, a single label is assigned to a bag of instances, addressing the problem we are discussing. In the past, MIL approaches typically involved pooling or attention mechanisms to fuse embeddings or setting up scoring functions  (Ilse et al., 2018; Feng and Zhou, 2017; Pinheiro and Collobert, 2015; Perini et al., 2023; Pang et al., 2023; Abati et al., 2019). In our case, we aim to transform this problem into a direct micro PU learning problem to achieve more accurate solutions and eliminate redundant information.
2.3. Anomaly Detection
@@
Anomaly Detection (AD) is referred to as the process of detecting data instances that significantly deviate from the majority of data instances  (Chandola et al., 2009; Pang et al., 2021). Historically, AD methods have been categorized into supervised AD  (Görnitz et al., 2013), weakly supervised AD  (Ruff et al., 2019; Pang et al., 2023; Perini et al., 2023), and unsupervised AD  (Ruff et al., 2018; Ye et al., 2023). However, there hasn’t been a specific AD algorithm tailored for the PU form, where only normal samples are labeled and there are no labeled anomalies. Therefore, existing methods cannot be used to address the proposed problem, as macro AD and macro PN classification are similarly ineffective due to high similarity.
3. Derivation of Unbiased and Balanced PU Learning
@@
In classical supervised learning problems, we often deal with positive-negative (PN) learning. Assume there is an underlying distribution $p(x,y)$, where $x\in\mathbb{R}^{d}$ is the input, and $y\in\{-1,+1\}$ is the output. Data of size $n_{+}$ are sampled from $p(x|y=+1)$, and data of size $n_{-}$ are sampled from $p(x|y=-1)$. We let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2}$ be a decision function from a function space $\mathcal{G}$, where $g_{-1}$ and $g_{+1}$ are the probabilities of the sample being negaftive and positive, respectively. We also let $\mathcal{L}:\mathbb{R}^{d}\times\{-1,+1\}\rightarrow\mathbb{R}$ be a loss function. The goal of PN learning is to use P data and N data to learn a classifier, denoted as $f:\mathbb{R}^{d}\rightarrow\{-1,+1\}$ which is based on the decision function $g$.
@@
(2) $E_{p(x)}[\mathcal{L}[g(x),-1]]=(1-\pi)E_{p(x|y=-1)}[\mathcal{L}[g(x),-1]]+\pi E_{p(x|y=+1)}[\mathcal{L}[g(x),-1]],$
@@
(3) $\hat{R}_{upu}(\hat{g})=\frac{\pi}{n_{p}}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),+1]+\frac{1}{n_{u}}\sum_{x_{i}\in U}\mathcal{L}[\hat{g}(x_{i}),-1]-\frac{\pi}{n_{p}}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),-1].$
Su et al.  (Su et al., 2021) argued that previous approaches struggle with balanced metrics. The objective of unbiased PU learning is to train a classifier that is unbiased when the class distribution of the test data matches that of the unlabeled data, rather than creating a balanced classifier. This can lead to poor performance for one of the classes, especially when $\pi$ is close to 0 or 1. In such cases, even if the model classifies all samples into a single class, achieving high accuracy, it does not meet the goals for real-world applications. Our objective is to learn a balanced classifier, despite the imbalance between positive and negative data in the unlabeled set. Building on the balanced PN learning objective, we aim to address this challenge:
(4) $R_{bpn}(g)=\frac{1}{2}E_{p(x|y=+1)}[\mathcal{L}[g(x),+1]]+\frac{1}{2}E_{p(x|y=-1)}[\mathcal{L}[g(x),-1]],$
@@
$\hat{R}_{bpu}(\hat{g})$ $=\frac{1}{2n_{p}}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),+1]+\frac{1}{2n_{u}(1-\pi)}\sum_{x_{i}\in U}\mathcal{L}[\hat{g}(x_{i}),-1]$
(5) $-\frac{\pi}{2n_{p}(1-\pi)}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),-1].$
@@
(6) $\hat{R}_{bpu}(g)=\frac{1}{2n_{p}(1-\pi)}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),+1]+\frac{1}{2n_{u}(1-\pi)}\sum_{x_{i}\in U}\mathcal{L}[\hat{g}(x_{i}),-1]-\frac{\pi}{2(1-\pi)}.$
@@
This simple formula illustrates a straightforward yet counterintuitive principle. When the loss function satisfies $\mathcal{L}[t,+1]+\mathcal{L}[t,-1]$ as a constant, directly treating unlabeled samples as negative and training the model using the expected loss for each class ̵supervisely results in a balanced learner. In other words, the model becomes unbiased with respect to the average accuracy metric. Building on this, we further derive the micro-level learning objective for PU learning.
Table 1. The table presents the AvgAcc of different algorithms under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$, along with the estimated AUC concerning changes in AvgAcc on the Imdb Dataset.
Method AvgAcc $AUC_{AvgAcc}$
2 3 4 5
Macro PN $76.75\pm 1.45$ $70.62\pm 0.83$ $62.61\pm 4.70$ $58.33\pm 4.81$ $67.08$
DeepSAD $52.17\pm 0.56$ $52.57\pm 1.20$ $55.01\pm 1.05$ $50.76\pm 0.77$ $52.63$
MIL ATT $82.58\pm 7.37$ $85.98\pm 1.32$ $80.66\pm 3.81$ $78.62\pm 1.54$ $81.96$
PReNET $84.53\pm 1.50$ $80.54\pm 1.50$ $71.78\pm 4.97$ $73.56\pm 2.49$ $77.60$
uPU $82.93\pm 1.14$ $76.30\pm 1.51$ $78.57\pm 5.37$ $72.29\pm 3.30$ $77.52$
nnPU $84.53\pm 1.49$ $58.39\pm 5.95$ $50.00\pm 0.00$ $50.00\pm 0.00$ $60.73$
BFGPU $88.13\pm 1.07$ $87.93\pm 0.58$ $86.02\pm 0.69$ $83.64\pm 0.78$ $86.43$

### v3

# Detecting Scarce and Sparse Anomalous: Solving Dual Imbalance in Multi-Instance Learning

@@
Abstract
In real-world applications, it is highly challenging to detect anomalous samples with extremely sparse anomalies, as they are highly similar to and thus easily confused with normal samples. Moreover, the number of anomalous samples is inherently scarce. This results in a dual imbalance Multi-Instance Learning (MIL) problem, manifesting at both the macro and micro levels. To address this "needle-in-a-haystack problem", we find that MIL problem can be reformulated as a fine-grained PU learning problem. This allows us to address the imbalance issue in an unbiased manner using micro-level balancing mechanisms. To this end, we propose a novel framework, Balanced Fine-Grained Positive-Unlabeled (BFGPU)-based on rigorous theoretical foundations. Extensive experiments on both synthetic and real-world datasets demonstrate the effectiveness of BFGPU.
00footnotetext: Corresponding Author: Yu-Feng Li and Lan-Zhe Guo.
1 Introduction
@@
In real-world applications, there is a strong demand for effective detection of anomalous samples, such as in quality inspection, risk control, and fault diagnosis  (Görnitz et al., 2013; Pang et al., 2023; Ye et al., 2023). Ideally, anomalous samples exhibit significant and easily distinguishable differences from normal samples, making them readily separable based on features or representations. However, in more realistic scenarios, the differences between anomalous and normal samples are often subtle because anomalies may manifest as sparse local information. As a result, anomalous samples are highly similar to normal ones at a macro level. Moreover, the number of anomalous samples is inherently scarce in reality, making them more difficult to identify.
On the one hand, this problem has significant practical relevance. For instance, in cancer cell detection, diseased samples are inherently rare, and within these samples, the pathological regions often occupy only a minuscule portion. This makes early detection before the cancer metastasizes exceptionally challenging. On the other hand, the rise of large language models (LLMs) has further amplified the importance of this detection problem. During the Reinforcement Learning from Human Feedback (RLHF) phase, aligning LLMs with human values requires training a separate detector as a reward model (Bai et al., 2022). This model is trained on human feedback to identify non-compliant sections in the LLM’s outputs. However, these non-compliant contents are both scarce and sparse. Compounding this is the high cost of human annotation, which typically only provides coarse-grained labels. These factors make training an accurate reward model extremely difficult, thereby further highlighting the critical nature of this detection challenge. An example is provided in fig.˜1, a real-world customer service quality inspection task, the goal is to detect instances of impoliteness within extended conversations between customer service and customer. Since the customer service is well-trained, mistakes—if any—typically occur in only a subtle utterance.
@@
Figure 1: This example illustrates the difficulty in distinguishing between normal and anomalous macro samples due to the low proportion of anomalous information within anomalous samples.
@@
Given the difficulty of distinguishing samples at the macro level, a promising approach is to seek solutions at the micro level—by decomposing each macro-level sample into multiple finer-grained components and performing discrimination at this finer resolution. However, there are no precise labels available at the micro level, so learning must rely solely on macro-level supervision. Multiple Instance Learning (MIL) is a paradigm of weakly supervised learning designed for scenarios with imprecise labels  (Waqas et al., 2024; Zhou, 2018; Gao et al., 2022). It structures data into two hierarchical levels: the macro level (referred to as bags) and the micro level (referred to as instances), and utilizes only macro-level labels for training. To address the lack of micro-level supervision, existing MIL methods often heuristically assign bag labels to all instances within the bag  (Ilse et al., 2018; Angelidis & Lapata, 2018; Pang et al., 2023). However, this strategy lacks theoretical grounding and can introduce substantial bias. In terms of data imbalance, the micro-level imbalance—known in MIL literature as the low witness rate problem—has been recognized but remains unresolved  (Carbonneau et al., 2018; Zhang et al., 2022). Meanwhile,macro-level imbalance has been largely overlooked and presents an even greater challenge.
@@
In this paper, instead of directly assigning bag labels to instances, we treat all instances from normal bags as positive instances and all instances from anomalous bags as unlabeled instances. This reformulates the MIL problem as a fine-grained Positive-Unlabeled (PU) learning task. PU learning is a learning paradigm that trains models using only positive and unlabeled data  (Bekker & Davis, 2020). Building upon this reformulation, we further derive a PU learning loss function based on rigorous theoretical analysis, which simultaneously addresses the dual imbalance challenge in MIL.
@@
Figure 2: This figure illustrates the different solution paradigms for the detection problem in this paper: coarse-grained positive-negative learning, MIL, and fine-grained positive-unlabeled learning.
In the text and image modalities, we conducted comprehensive comparisons with existing supervised learning, anomaly detection, MIL, and PU learning methods based on customer service quality inspection (CSQI) and invasive ductal carcinoma (IDC) tasks, respectively. Furthermore, to validate the robustness of our algorithm under varying macro and micro imbalance ratios, we synthesized numerous datasets with different imbalance ratios based on four fundamental sentiment analysis datasets for additional comparative experiments. We also compared its performance with popular LLMs in addressing the "needle in a haystack" problem (Wang et al., 2024a; Kuratov et al., 2024). Finally, through ablation studies and parameter sensitivity analysis, we further enriched the discussion on the algorithm’s robustness.
@@
Our Contributions. 1. We formalize the challenging detection problem in practical applications and propose a dual imbalanced MIL problem. 2. We propose a novel perspective that reformulates the MIL problem as a balanced PU learning task, which addresses the issues of bias and imbalance. 3. We incorporate macro information into the learning process, leading to BFGPU, an algorithm that enables balanced and unbiased macro performance by training at the micro level. 4. We provide a theoretical analysis demonstrating the advantages of BFGPU over macro-level learning and conventional MIL approaches. 5. Extensive experiments validate the superior performance of BFGPU.
2 Related Works
@@
PU learning can utilize positive data and unlabeled data to train a classifier that distinguishes between positive and negative data  (Bekker & Davis, 2020). Most PU learning algorithms aim for the model to be unbiased in terms of accuracy, and there are many effective algorithms currently available  (Du Plessis et al., 2015; Kiryo et al., 2017; Chen et al., 2020; Kato et al., 2018; Shi et al., 2018; Sansone et al., 2018; Hou et al., 2018; Hsieh et al., 2019, 2015; Wang et al., 2024b). Some PU learning algorithms address the issue of imbalance in the positive-negative ratio within unlabeled data and strive for unbiased average accuracy (AvgAcc)  (Su et al., 2021). Additionally, theoretical research on PU learning has demonstrated its superiority  (Niu et al., 2016).
@@
Multi-instance learning (MIL) is a form of weakly supervised learning  (Zhou, 2018) characterized by inexact supervision, where labels are not precise  (Carbonneau et al., 2018). In MIL, a single label is assigned to a bag of instances, addressing the problem we are discussing. In the past, MIL approaches typically involved pooling or attention mechanisms to fuse embeddings or setting up scoring functions  (Ilse et al., 2018; Feng & Zhou, 2017; Pinheiro & Collobert, 2015; Perini et al., 2023; Pang et al., 2023; Abati et al., 2019). In our case, we aim to transform this problem into a direct micro PU learning problem to achieve more accurate solutions and eliminate redundant information.
@@
Anomaly Detection (AD) the process of detecting data instances that significantly deviate from the majority of data instances  (Pang et al., 2021). Historically, AD methods have been categorized into supervised AD  (Görnitz et al., 2013), weakly supervised AD  (Ruff et al., 2019; Pang et al., 2023; Perini et al., 2023), and unsupervised AD  (Ruff et al., 2018; Ye et al., 2023). However, there hasn’t been an effective AD algorithm tailored for the form where anomalous are both scarce and sparse.
3 Derivation of Unbiased and Balanced PU Learning
@@
In classical supervised learning problems, we often deal with positive-negative (PN) learning. Assume there is an underlying distribution $p(x,y)$, where $x\in\mathbb{R}^{d}$ is the input, and $y\in\{-1,+1\}$ is the output. Data of size $n_{+}$ are sampled from $p(x|y=+1)$, and data of size $n_{-}$ are sampled from p(x|y=1)p(x|y=-1). We let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2}$ be a decision function from a function space $\mathcal{G}$, where $g_{-1}$ and $g_{+1}$ are the probabilities of the sample being negaftive and positive, respectively. We also let $\mathcal{L}:\mathbb{R}^{d}\times\{-1,+1\}\rightarrow\mathbb{R}$ be a loss function. The goal of PN learning is to use P data and N data to learn a classifier, denoted as $f:\mathbb{R}^{d}\rightarrow\{-1,+1\}$ which is based on the decision function $g$.
@@
$E_{p(x)}[\mathcal{L}[g(x),-1]]=(1-\pi)E_{p(x|y=-1)}[\mathcal{L}[g(x),-1]]+\pi E_{p(x|y=+1)}[\mathcal{L}[g(x),-1]],$ (2)
@@
$\hat{R}_{upu}(\hat{g})=\frac{\pi}{n_{p}}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),+1]+\frac{1}{n_{u}}\sum_{x_{i}\in U}\mathcal{L}[\hat{g}(x_{i}),-1]-\frac{\pi}{n_{p}}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),-1].$ (3)
(Su et al., 2021) argued that previous approaches struggle with balanced metrics. The objective of unbiased PU learning is to train a classifier that is unbiased when the class distribution of the test data matches that of the unlabeled data, rather than creating a balanced classifier. This can lead to poor performance for one of the classes, especially when $\pi$ is close to 0 or 1. In such cases, even if the model classifies all samples into a single class, achieving high accuracy, it does not meet the goals for real-world applications. Our objective is to learn a balanced classifier, despite the imbalance between positive and negative data in the unlabeled set. Building on the balanced PN learning objective, we aim to address this challenge:
$R_{bpn}(g)=\frac{1}{2}E_{p(x|y=+1)}[\mathcal{L}[g(x),+1]]+\frac{1}{2}E_{p(x|y=-1)}[\mathcal{L}[g(x),-1]],$ (4)
@@
[…]
$-\frac{\pi}{2n_{p}(1-\pi)}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),-1].$ (5)
@@
$\hat{R}_{bpu}(g)=\frac{1}{2n_{p}(1-\pi)}\sum_{x_{i}\in P}\mathcal{L}[\hat{g}(x_{i}),+1]+\frac{1}{2n_{u}(1-\pi)}\sum_{x_{i}\in U}\mathcal{L}[\hat{g}(x_{i}),-1]-\frac{\pi}{2(1-\pi)}.$ (6)
@@
This simple formula illustrates a straightforward yet counterintuitive principle. When the loss function satisfies $\mathcal{L}[t,+1]+\mathcal{L}[t,-1]$ as a constant, directly treating unlabeled samples as negative and training the model using the expected loss for each class␣supervisely results in a balanced learner. In other words, the model becomes unbiased with respect to the average accuracy metric. Building on this, we further derive the micro-level learning objective for PU learning.
4 Balanced Fine-Grained PU Learning
4.1 Micro-to-Macro Optimization
[…]
$G([g(x_{i1}),g(x_{i2}),\cdots,g(x_{il})])=\sum_{j=1}^{l}p(j=\arg\max_{k}g_{-1}(x_{ik}))g(x_{ij})$ (8)
Referring to eq.˜8, we aim to find the idealized $p$ which makes $p(j=\arg\max_{k}g_{-1}(x_{ik}))=\mathbb{I}[j=\arg\max_{k}g_{-1}(x_{ik})]$. However, since we cannot obtain $g$, during the training process of neural networks, $\hat{g}$ from the neural network often deviates somewhat from the ideal decision function $g$. Consequently, we cannot directly obtain the value of the indicator function $\mathbb{I}[j=\arg\min_{k}g_{-1}(x_{ik})]$, but we can estimate the probability of the condition being satisfied. So, we use the probability function $\hat{p}$ instead of the function $p$.
$\hat{p}(x_{ij})=\frac{\exp(\hat{g}_{-1}(x_{ij}))}{\sum_{k=1}^{l}\exp(\hat{g}_{-1}(x_{ik}))}$ (9)
[…]
$+$ $\frac{1}{2\cdot|N_{macro}|\cdot l}\sum_{X_{i}\in N_{macro}}\mathcal{L}[G([\hat{g}(x_{i1}),\cdots,\hat{g}(x_{il})]),-1]$ (10)
Based on eqs.˜6, 8, 9 and 4.1, through the transformation of the learning paradigm, we derive a loss function that enables unbiased and balanced optimization of macro-level performance directly at the micro level which is called balanced fine-grained PU learning loss:
[…]
$+$ $\frac{1}{2|U_{micro}|}\sum_{x_{ij}\in U_{micro}}\hat{p}(x_{ij})\mathcal{L}[\hat{g}(x_{ij}),-1]$ (11)
4.2 Pseudo Labels Based on Macro Information
$N_{pse}=\{(x_{ij},-1),j=\arg\max_{k}\hat{g}_{-1}(x_{ik}),X_{i}\in N_{macro}\}$ (12)
$P_{pse}=\{(x_{ij},+1),j=\arg\min_{k}\hat{g}_{-1}(x_{ik}),X_{i}\in N_{macro}\}$ (13)
$\hat{R}_{pse}(\hat{g})=\sum_{(x_{i},y_{i})\in N_{pse}+P_{pse}}\mathcal{L}[\hat{g}(x_{i}),y_{i}]$ (14)
4.3 Adjusted Decision Threshold (ADT)
$T=sort([\hat{g}_{-1}(x_{i}),x_{i}\in U_{micro}])[\lfloor|U_{micro}|\cdot\pi\rfloor]$ (15)
This approach helps determine an accurate threshold because the decision function $g(x)$ reflects the conditional probability $p(y|x)$. $p(y|x)$ remains consistent between training and testing data. Therefore, the threshold obtained on the unlabeled data is highly applicable to the test data. The overall process of the algorithm can be seen in algorithm˜1.
5 Experiments
5.1 Experimental Settings
For textual data, we primarily employ RoBERTa (Liu et al., 2019) as the backbone model (except for ablation studies concerning the backbone itself). For image data, we utilize ResNet-18 as the backbone to ensure a fair comparison across all algorithms. We uniformly adopt Adam as the optimizer  (Kingma & Ba, 2015), with a learning rate set to $10^{-5}$, and CosineAnnealingLR as the learning rate scheduler. The number of epochs was set to $5$, batch size to $16$. We denote $\sigma_{micro}$ as the imbalance ratio at the micro level, which represents the positive-negative ratio of micro samples in negative macro samples. $\sigma_{macro}$ represents the imbalance ratio at the macro level, which is the ratio of positive macro samples to negative macro samples. We set $\lambda_{bfgpu}=1/\pi$ and $\lambda_{pse}=1$ in all experiments. For micro-level PU learning, $\pi=\frac{\sigma_{micro}}{\sigma_{micro}+1}$. The algorithms were implemented using the PyTorch framework  (Paszke et al., 2019). We used average accuracy and F1 score, two commonly used evaluation metrics in imbalanced learning. Regarding the F1 score, when dealing with class imbalance problems, we typically focus on the performance of the minority (anomalous) class. We also compared the algorithm performance as $\sigma_{micro}$ varies, using the area under the curve $AUC_{AvgAcc}$ and $AUC_{F1}$. 4 A800 GPUs are used for all experiments.
5.2 Compared Methods
Our comparative algorithms consist of: Macro Supervised: Conventional macro binary classification. Macro DeepSAD: Supervised macro AD using DeepSAD (Ruff et al., 2019). Macro Imbalanced Learning: Under sampling and Over sampling. MIL: MIL algorithms used to address inexact supervision, using five multi-instance learning algorithms MIL-MaxPooling, MIL-TopkPooling (k was set to 3), MIL-Attention  (Ilse et al., 2018), MIL-FGSA  (Angelidis & Lapata, 2018), and MIL-PReNET (Pang et al., 2023). Micro DeepSAD: Supervised micro AD, using DeepSAD  (Ruff et al., 2019) and treating unlabeled samples as negative class. Micro DeepSVDD: Unsupervised micro AD, using DeepSVDD  (Ruff et al., 2018) which is an One-Class Classification (OCC) method using only positive samples. Micro PU Learning: Utilizing four types of loss functions uPU  (Du Plessis et al., 2015), nnPU  (Kiryo et al., 2017), balancedPU  (Su et al., 2021), and robustPU (Zhu et al., 2023) for micro-level PU learning. LLMs: GPT-4-turbo (Achiam et al., 2023), Qwen3 (Yang et al., 2025), Deepseek v3 (Liu et al., 2024), Gemini2.5 (Comanici et al., 2025).
Table 1: The comparison on the IDC dataset.
Method AvgAcc F1 Score
Macro Supervised 58.70 57.13
DeepSAD 44.30 43.20
MIL Maxpooling 65.55 62.39
Topkpooling 63.74 63.82
Attention 56.03 49.10
FGSA 60.22 59.86
PReNET 49.30 38.46
Micro DeepSAD 61.94 58.81
DeepSVDD 52.15 49.10
uPU 56.88 56.36
nnPU 62.66 61.00
balancePU 57.33 55.73
robustPU 64.38 63.16
BFGPU 73.87 73.21
Table 2: The comparison on the CSQI dataset.
Method AvgAcc F1 Score
Macro Supervised $52.38$ $8.33$
DeepSAD $47.61$ $6.67$
MIL Maxpooling $48.81$ $46.67$
Topkpooling $52.38$ $45.45$
Attention $54.76$ $53.27$
FGSA $53.57$ $54.14$
PReNET $59.52$ $59.05$
Micro DeepSAD $58.33$ $63.12$
DeepSVDD $50.00$ $0.00$
uPU $50.00$ $0.00$
nnPU $50.00$ $0.00$
balancedPU $50.00$ $0.00$
robustPU $50.00$ $0.00$
BFGPU $71.43$ $75.31$
5.3 Experiments with the IDC Dataset
We evaluated various methods with a standard MIL dataset IDC, which is a medical image classification dataset primarily used for breast cancer detection research (Bolhasani et al., 2020). This dataset contains patches of breast tissue images labeled according to the presence of invasive ductal carcinoma. In the dataset setup, each bag is a patient and the $\sigma_{macro}$ is around $3$. Each instance is a 50×50 pixel image patch, and each bag contains 4 instances. The results are shown in table˜1.
5.4 Experiments with the CSQI Dataset
We evaluated various methods with a real-world CSQI dataset depicted in fig.˜1. The dataset consists of instances where service personnel were flagged for substandard performance during service. In CSQI, anomalous samples are extremely scarce, with only 24 dialogue sessions, while normal samples are abundant. we set $\sigma_{macro}=100$ and randomly sampled 2,400 normal sessions for the experiments. The $\sigma_{micro}$ varied inconsistently, averaging around $25$. The results are shown in table˜2.
5.5 Experiments with Synthetic Datasets
Since real datasets cannot control the imbalance ratios $\sigma_{macro}$ and $\sigma_{micro}$, we use sentiment analysis datasets to synthesize datasets with different imbalance ratios to explore the capability of various algorithms in handling dual imbalance problems. For example, we randomly combine 5 positive samples and 1 negative sample from the original dataset to synthesize a positive sample with $\sigma_{micro}=5$. For short-text ones SST-2  (Socher et al., 2013) and Sentiment140  (Go et al., 2009), $\sigma_{micro}$ was set to $[2,4,6,8,10]$. For long-text datasets IMDB (Maas et al., 2011) and Amazon, $\sigma_{micro}$ was set to $[2,3,4,5]$, the remaining experimental results can be found in. All experiments were repeated $3$ times using random seeds $[0,1,2]$. The results can be found in tables˜3, 5, 7, 6 and 8.